# Nonlocal characteristics of two-qubit gates and their argand diagrams

### v1

@@ IV Pair of orthonormal product states to maximally entangled states
To show that there exist perfect entanglers without any chord passing through origin but, capable of transforming a pair of orthonormal product states into maximally entangled states, we consider the triangular region with vertices $(\pi/4,\pi/4,\pm\pi/4)$ and $(\pi/2,\pi/2,0)$ in the Weyl chamber [FIG. 3a]. It is a region of perfect entanglers with $h_{1}=h_{4}$. For perfect entanglers represented by the edges of the triangle at least one chord passes through origin and for those represented by the interior points only the conditions given in cases 1 and 2, discussed above, are applicable. It can be verified that the conditions $(|\phi_{1}|^{2}=0.5,|\phi_{2}|^{2}=0.5,~{}\text{and}~{}|\phi_{3}|^{2}=0.5)$ described in case 1 are not satisfied by any perfect entanglers represented by the interior points and only the line connecting $\sqrt{\text{iSWAP}}$ and iSWAP [FIG. 3a] satisfy the condition $(|\phi_{2}|^{2}=|\phi_{3}|^{2})$ described in case 2. Thus, in this perfect entanglers region without any chord passing through origin, only the perfect entanglers represented by the violet colored line (shown in FIG. 3) can transform a pair of orthonormal product states into maximally entangled states and the perfect entanglers represented by other interior points cannot do such a transformation.
@@ V Quantification of nonlocal characteristics using chords
Like two-qubit states, two-qubit gates also possess entanglement and it can be quantified using linear entropy [6, 18]. Two-qubit gates represented by $c_{3}$ axis of Weyl chamber have maximum value of linear entropy [19]. Entangling power is another nonlocal measure of two-qubit gates which quantifies the average entanglement generated over uniform distribution of product states [8, 9]. Gate typicality is a nonlocal measure which is complementary to entangling power [11, 12]. Both entangling power $(e_{p})$ and gate typicality $(g_{t})$ together describe the linear entropy $(L)$ of two-qubit gates. The relationship between them can be written as follows.
$L=\dfrac{3}{8}\left[3e_{p}+g_{t}\right].$ (21)
Recently, entangling power of two-qubit gates was shown to be proportional to mean squared length of the chords in the argand diagram of two-qubit gates [13]. Entangling power can be written in terms of chord lengths as follows.
$e_{p}=\dfrac{1}{72}\sum_{j=1}^{3}\sum_{k>j}|e^{ih_{j}}-e^{ih_{k}}|^{2}.$ (22)
@@
$e_{p}(U)=\dfrac{4}{9}\left(L(U)+\left[L(US)-L(S)\right]\right),$ (23)
@@
$g_{t}(U)=\dfrac{4}{3}\left(L(U)-\left[L(US)-L(S)\right]\right).$ (24)
@@
$g_{t}=\dfrac{1}{12}\sum_{j=1}^{3}\sum_{k>j}|e^{ih_{j}}-e^{-ih_{k}}|^{2}.$ (25)
This replacement of $e^{ih_{k}}$ by its complex conjugate in Eq. 22 can be geometrically interpreted as follows: All the three points $e^{ih_{j}}$, $e^{ih_{k}}$, and $e^{-ih_{k}}$ in the argand diagram define a triangle as shown in FIG. 4. The side connecting $e^{ih_{k}}$ and $e^{-ih_{k}}$ is parallel to the imaginary axis. The side connecting $e^{ih_{j}}$ and $e^{ih_{k}}$, and the side connecting $e^{ih_{j}}$ and $e^{-ih_{k}}$ are reflection of each other about the median from $e^{ih_{j}}$. We refer to the chord connecting $e^{ih_{j}}$ and $e^{-ih_{k}}$ as the median reflected chord corresponding to the chord connecting $e^{ih_{j}}$ and $e^{ih_{k}}$. For each chord defining entangling power (shown in FIG. 2a) their corresponding median reflected chords define gate typicality. Thus, replacing the symmetric combination of $L(U)$ and $L(US)-L(S)$ in the definition of entangling power by its antisymmetric combination to define gate typicality corresponds to replacing the squared length of the chords describing the entangling power by the squared length of their corresponding median reflected chords. It has to be noted that if any chord describing the entangling power is parallel to imaginary axis, then the length of the corresponding chord describing the gate typicality is zero.
@@
$g_{t}=\dfrac{1}{6}\sum_{k\neq j}|e^{ih_{j}}-e^{-ih_{k}}|^{2},~{}~{}\text{for}~{}\text{any}~{}j.$ (26)
@@
Since both entangling power and gate typicality of a two-qubit gate can be calculated using the chords present in their argand diagram, the linear entropy (Eq. 21) describing the operator entanglement of the two-qubit gate can also be calculated using the chords as follows.
$L=\dfrac{1}{64}\sum_{j=1}^{3}\sum_{k>j}\left[|e^{ih_{j}}-e^{ih_{k}}|^{2}+2|e^{ih_{j}}-e^{-ih_{k}}|^{2}\right]$ (27)
@@ VI Conclusion
To conclude, the argand diagram of squared eigenvalues of nonlocal part of two-qubit gates are very useful to calculate the nonlocal measures such as entangling power, gate typicality, and operator entanglement of two-qubit gates. In the argand diagram of a two-qubit gate, for each chord describing the entangling power, there exist a median reflected chord describing the gate typicality. We have identified the planes containing the perfect entanglers with chords passing through origin. Among SPEs, only the points representing CNOT and iSWAP equivalence classes are present in more than two planes and all other SPEs are contained in two planes. $\sqrt{\text{SWAP}}$ and $\sqrt{\text{SWAP}}^{\dagger}$ equivalence classes are the only perfect entanglers contained in three planes. Perfect entanglers with a chord passing through origin can transform a pair of orthonormal product states into maximally entangled states. However, not all the perfect entanglers without any chord passing through origin can transform a pair of orthonormal product states into maximally entangled states; only those satisfying one of the conditions derived in section IV can do such a transformation.

### v2

@@ IV Perfect entanglers without any chord passing through origin
Case 2 describes the perfect entanglers on $c_{1}=c_{2}$ and $c_{2}=\pm c_{3}$ planes of Weyl chamber. On $c_{1}=c_{2}$ plane, the perfect entanglers are contained in the triangular region with vertices $[\pi/4,\pi/4,\pm\pi/4]$ and $[\pi/2,\pi/2,0]$ [Fig. 3a]. All three edges of the triangular region represent perfect entanglers with at least a chord passing through origin. The interior is a region of perfect entanglers with $h_{1}=h_{4}$ and hence only the condition $|\phi_{2}|^{2}=|\phi_{3}|^{2}$ is applicable. It can be verified that only $U_{d}(c_{1},c_{1},0)$ with $c_{1}\in(\pi/4,\pi/2)$ satisfy the condition $|\phi_{2}|^{2}=|\phi_{3}|^{2}$. Thus, in this region of perfect entanglers without any chord passing through origin, only the perfect entanglers represented by the open line segment between the points $[\pi/4,\pi/4,0]$ and $[\pi/2,\pi/2,0]$ (violet coloured line in Fig. 3a) can transform a pair of orthonormal product states into maximally entangled states and the perfect entanglers represented by other interior points cannot do such a transformation. Similarly, it can be shown that on $c_{2}=\pm c_{3}$ planes only $U_{d}(\pi/2,c_{2},\pm c_{2})$ with $c_{2}\in(0,\pi/4)$ (purple and orange coloured lines in Fig. 3a) can transform a pair of orthonormal product states into maximally entangled states. Conditions obtained in case 1 are applicable to the perfect entanglers represented by the points other than those existing on the five planes discussed in the previous section and $c_{1}=c_{2}$, $c_{2}=\pm c_{3}$ planes of perfect entanglers polyhedron [Fig. 1].
@@ V Quantification of nonlocal characteristics using chords
Like two-qubit states, two-qubit gates also possess entanglement and can be quantified using linear entropy [6, 20]. Two-qubit gates represented by $c_{z}$ axis of the Weyl chamber have maximum value of linear entropy [21]. Entangling power is another nonlocal measure of two-qubit gates that quantifies the average entanglement generated over uniform distribution of product states [8, 9]. Gate typicality is a nonlocal measure which is complementary to entangling power [11, 12]. Both entangling power $(e_{p})$ and gate typicality $(g_{t})$ together describe the linear entropy $(L)$ of two-qubit gates. The relationship between them can be written as follows.
$L=\frac{3}{8}\left[3e_{p}+g_{t}\right].$ (41)
Recently, the entangling power of two-qubit gates was shown to be proportional to the mean squared length of the chords in the argand diagram of two-qubit gates [13]. Entangling power can be written in terms of squared chord lengths as follows.
$e_{p}=\frac{1}{72}\sum_{j=1}^{3}\sum_{k>j}|e^{ih_{j}}-e^{ih_{k}}|^{2}.$ (42)
@@
$e_{p}(U)=\frac{4}{9}\left(L(U)+\left[L(US)-L(S)\right]\right),$ (43)
@@
$g_{t}(U)=\frac{4}{3}\left(L(U)-\left[L(US)-L(S)\right]\right).$ (44)
@@
$g_{t}=\frac{1}{12}\sum_{j=1}^{3}\sum_{k>j}|e^{ih_{j}}-e^{-ih_{k}}|^{2}.$ (45)
This replacement of $e^{ih_{k}}$ by its complex conjugate in Eq. 42 can be geometrically interpreted as follows: All the three points $e^{ih_{j}}$, $e^{ih_{k}}$, and $e^{-ih_{k}}$ in the argand diagram define a triangle as shown in Fig. 4. The side connecting $e^{ih_{k}}$ and $e^{-ih_{k}}$ is parallel to the imaginary axis. The side connecting $e^{ih_{j}}$ and $e^{ih_{k}}$, and the side connecting $e^{ih_{j}}$ and $e^{-ih_{k}}$ are reflection of each other about the median from $e^{ih_{j}}$. We refer to the chord connecting $e^{ih_{j}}$ and $e^{-ih_{k}}$ as the median reflected chord corresponding to the chord connecting $e^{ih_{j}}$ and $e^{ih_{k}}$. For each chord defining entangling power (shown in Fig. 2a) their corresponding median reflected chords define gate typicality. Thus, replacing the symmetric combination of $L(U)$ and $L(US)-L(S)$ in the definition of entangling power by its antisymmetric combination to define gate typicality corresponds to replacing the squared length of the chords describing the entangling power by the squared length of their corresponding median reflected chords. It has to be noted that if any chord describing the entangling power is parallel to the imaginary axis, then the length of the corresponding chord describing the gate typicality is zero.
@@
$g_{t}=\frac{1}{6}\sum_{k\neq j}|e^{ih_{j}}-e^{-ih_{k}}|^{2},~{}~{}\textnormal{for}~{}\textnormal{any}~{}j.$ (46)
@@
Since both the entangling power and gate typicality of a two-qubit gate can be calculated using the chords present in their argand diagram, the linear entropy (Eq. 41) describing the operator entanglement of the two-qubit gate can also be calculated using the chords as follows.
$L=\frac{1}{64}\sum_{j=1}^{3}\sum_{k>j}\left[|e^{ih_{j}}-e^{ih_{k}}|^{2}+2|e^{ih_{j}}-e^{-ih_{k}}|^{2}\right]$ (47)
@@ VI Conclusion
To conclude, the argand diagram of squared eigenvalues of nonlocal part of two-qubit gates are very useful to calculate the nonlocal measures such as entangling power, gate typicality, and operator entanglement of two-qubit gates. In the argand diagram of a two-qubit gate, for each chord describing the entangling power, there exist a median reflected chord describing the gate typicality. We have identified the planes containing the perfect entanglers with chords passing through origin. Among SPEs, only the points representing CNOT and iSWAP equivalence classes are present in more than two planes and all other SPEs are contained in two planes. $\sqrt{\textnormal{SWAP}}$ and $\sqrt{\textnormal{SWAP}}^{\dagger}$ equivalence classes are the only perfect entanglers contained in three planes. Perfect entanglers with a chord passing through origin can transform a pair of orthonormal product states into maximally entangled states. However, not all the perfect entanglers without any chord passing through origin can transform a pair of orthonormal product states into maximally entangled states; only those satisfying one of the conditions derived in this paper can do such a transformation.